OpenStudy (1645323):

A box is made by cutting out squares of side x from an 8 in by 12 in piece of cardboard. The polynomial function V(x) = (8 - 2x)(12 - 2x)(x) represents the volume of box for any size square x. Multiply this expression to find the standard form for this polynomial function. Select one: a. V(x) = 4x3 - 32x2 + 96x b. V(x) = -4x3 + 96x c. V(x) = -x3 - 96 d. V(x) = x3 - 32x2 + 96x

OpenStudy (1645323):

@hartnn

OpenStudy (ddcamp):

FOIL the first two terms, then multiply everything by x.

OpenStudy (geekfromthefutur):

Ok im going to help you ok

OpenStudy (geekfromthefutur):

???

OpenStudy (1645323):

so 96+-16x+-24x+4x^2

OpenStudy (geekfromthefutur):

EXAMPLE V = 144x - 52x^2 + 4x^3 dV/dx = 144 - 104x + 12x^2 To maximize volume, set the derivative to zero: 12x^2 - 104x + 144 = 0 x = (-(-104) +/- sqrt((-104)^2 - 4(12)(144))) / (2*12) x = (104 +/- sqrt(10816 - 6912)) / 24 x = (104 +/- sqrt(3904)) / 24 x =~ 6.9 or 1.7 But x can't be more than 4, or the length of the box would be negative, so therefore: Height: x =~ 1.7 Length: 8 - 2x =~ 4.54 Width: 18 - 2x =~ 14.54 V =~ 114.2

OpenStudy (1645323):

4x^3-40x^2+96x

OpenStudy (ddcamp):

Yup!

OpenStudy (1645323):

is that the answer?

OpenStudy (ddcamp):

That seems right, but it's not one of the answers...

OpenStudy (1645323):

ok thanks